Sam wants to build a fence in a rectangular space in his back yard. This will be a safe place for the children to play. He has 40 meters of fencing. Assume that he is going to use all the fencing. If the width of the rectangular fenced area is 8 meters, what would the length of the rectangular area be? How many square meters would Bill's children have to play in?
Answer by wrex(1)

40=2(L+8)
L+8=20
L=12
Area=L*W
Area=12*8= 96 sq.ft.